OpenStudy (ladylauren):

A system of two equations is shown below: Equation C: a = 2b + 7 Equation D: a = 5b − 3 What value of a can be substituted into equation D to solve the system of equations? 2b 5b 5b − 3 2b + 7

OpenStudy (ladylauren):

i think its 2b +7

OpenStudy (princeharryyy):

yes right u are. :)

OpenStudy (ladylauren):

can you help me with this? Which set of steps shows the solution to the equation 4y = −12? y = −12 − 4; y = 8 y = −12 ÷ (−4); y = 3 y = −12 − (−4); y = −8 y = −12 ÷ 4; y = −3

OpenStudy (princeharryyy):

y = −12 ÷ 4; y = −3, u just divide by 4 on both sides.

OpenStudy (ladylauren):

so B?

OpenStudy (princeharryyy):

D

OpenStudy (ladylauren):

ohh ok

OpenStudy (princeharryyy):

:)

OpenStudy (ladylauren):

thx!!

OpenStudy (princeharryyy):

anything else?

OpenStudy (ladylauren):

no but thx.

OpenStudy (princeharryyy):

okay. i gotta go sleep than, it's already 3am.

OpenStudy (ladylauren):

wow its 5:29pm here

OpenStudy (princeharryyy):

different countries different time zones. :)

OpenStudy (ladylauren):

yup

OpenStudy (princeharryyy):

:) bye. tc